Find the length and width of a rectangle whose perimeter is 40 feet, and whose area is 96 square feet.
I think I have the right equation for the perimeter: P=2l+2w=96 sq ft.
Is this right?
I do not know how to set the other equation up or solve it.
thanks

Answer by stanbon(75887)

Find the length and width of a rectangle whose perimeter is 40 feet, and whose area is 96 square feet.
I think I have the right equation for the perimeter: P=2l+2w=40 ft.
----------
Then, divide thru by 2 to get:
l + w = 20
-----------
If width = w, length = 20-w
----------------
AREA:
Area = (length)(width)
96 = w(20-w)
96 = 20w-w^2
-----
w^2 -20w+96 = 0
(w-8)(w-12)=0
If w=8 then length = 12
If w=12 then length = 8
==================
Cheers,
Stan H.



I do not know how to set the other equation up or solve it.